The sum of the measures of a complement and supplement of an angle is 200. What is the measure of the angle?
Answer by rapaljer(4667)

Let x=the angle
90-x=the compliment
180-x=the supplement

90-x + 180-x = 200
270 - 2x =200
-2x = -70
x=35 The angle
90-x =55 The Complement
180-x = 145 The Supplement

Check: 55 + 145 = 200 It checks!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida